OpenStudy (anonymous):

Find the linearization of the function given below at the specified point. f(x) = 2 ln(x) at a = 1

myininaya (myininaya):

f'(x)=2*1/x f'(1)=2*1/1=2 is slope we know a point on this line (1,f(1)=2ln(1)=2(0)=0)=(1,0) y=mx+b 0=2*1+b -2=b y=2x-2 L(x)=2x-2